OpenStudy (destinykiara99):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1. Show your work for full credit

satellite73 (satellite73):

\[\theta =-\frac{3\pi}{4}\]?

satellite73 (satellite73):

you got a unit circle cheat sheet?

satellite73 (satellite73):

you do now

OpenStudy (destinykiara99):

Yes, that is the equation :)

satellite73 (satellite73):

take a look at the last page, the unit circle, and locate \(-\frac{3\pi}{4}\)

satellite73 (satellite73):

you see the x and y coordinates of that point?

satellite73 (satellite73):

you have to go around the unit circle clockwise 

satellite73 (satellite73):

is it clear what the coordinates of that point are on the unit circle?

OpenStudy (destinykiara99):

No, I'm sorry this whole lesson I'm having a hard time with

satellite73 (satellite73):

it is the same as the point labelled \(\frac{5\pi}{4}\)

satellite73 (satellite73):

in any case you see that both the x and y are the same there

satellite73 (satellite73):

so if the x is -1, so is the y

OpenStudy (destinykiara99):

Oh, okay